Known Answer: W12=-51.0 Btu, constant Cv: Q12= -19.2 Btu, Table A-22E: Q12=-18.8 Btu....... Please explain in detail how to use the table.3.145 Air is compressed in a piston-cylinder assembly from pi = 10 lbf/in?, T1 = 500°R to a final volume of K2 = 1 ft in a prTable A-22E Ideal Gas Properties of Air | T (R), h and u (Btu/Ibm), sº (Btu/lbm•°R) | when As = 0 ThusºpV, 360 85.97 61.29 0Table A-22E Ideal Gas Properties of Air (continued) I (R), h and u (Btu/lbm), sº (Btu/lbmºR) T (R), h and u (Btu/lbm), sº (Bt
Answer #1

The detailed solution is given below:

Goven- PE 10 dbf lim? T= 500R M = 0.5 lbm = 183 I n = 125 boly tro bic brocess # - [tral take paran > B = 9(0- ne )_0 вч, уг

# work donl, Wi-a = pdv = Pz Vz - PV -n (16153 psi Jx (1f4?) – (lo ksi) x (9.26 ft ?) - 1.25 = 161.53— 92.6 -0.25 = -275.72 p

# The energy bakancel: O-W = AU [sincl K.E=P.E=0] => Q = wtau (al At T- 500R ZAOF Cv = 0.171 Brullbmo R [from ideal gas Speci

=> Q12 = (-51.02)+ (0.5)x (149.61 – 85.20) = -18.81 644 Interpolation: U, -Venik = levici (bero-berada) que = 14150 43983-155